Work done during reversible isothermal process is given by

Grade:11

1 Answers

shekharnandanwar
6 years ago
In isothermal process temperature is constant the reversible work involved when a gas changes from state A to state B is{\displaystyle W_{A\to B}=-\int _{V_{A}}^{V_{B}}p\,dV} W_{{A\to B}}=-\int _{{V_{A}}}^{{V_{B}}}p\,dV In reverse process this work done is equals the area under the relevant pressure-volume isotherm.